OpenStudy (anonymous):

remind me: Is there an easier way to solve this i^132

OpenStudy (anonymous):

take the integer remainder when you divide \(132\) by \(4\)

OpenStudy (anonymous):

is i a complex number ? u mean to say that

OpenStudy (anonymous):

since 4 divides 132 evenly, the remainder is zero that means \[i^{132}=e^0=1\]

OpenStudy (anonymous):

yes and so if say for another example was i^75 . . . all i have to do is to divide by 4?

OpenStudy (anonymous):

when i say "divide' i mean take the whole number remainder in the case of \(75\) it is \(3\) because 4 goes in to 72 evenly, leaving a remainder of 3

OpenStudy (anonymous):

that makes \[i^{75}=i^3=-i\]

OpenStudy (anonymous):

i^4n =1 i^4n+1 =i i^4n +2 =-1 i^4n+3= -i (where n belongs to natural numbers) Just remember this

OpenStudy (anonymous):

that patter is \[i^1=i\\ i^2=-1\\ i^3=-i\\ i^4=1\] and so on

OpenStudy (anonymous):

ok i understand the dividing part but why isnt it 1. How was it -i

OpenStudy (anonymous):

as in why its i^4=1.

OpenStudy (anonymous):

What is i is is basically \[\sqrt{1}\] square of that becomes -1 and square of -1 becomes 1 so i^4 =1

OpenStudy (anonymous):

These values that i @satellite73 wrote are proved accordingly

OpenStudy (anonymous):

alright thanks got it